OpenStudy (dallaslifebaby01):

Write a linear equation in Standard form. Tell how you know it is linear. (5 pts)

OpenStudy (tkhunny):

You should give that a shot. y = mx+b -- slope-intercept form x/a + y/b = 1 -- Intercept form (y-y1) = m(x-x0) -- Point-slope form Where's the Standard Form? Is it different from Normal Form?